OpenStudy (anonymous):

The equation, y = 10(1.0605)^x , represents the winning prize money, in thousands, for first place in the Indianapolis 500 auto race. x represents the number of years after 1911, the first year. What is the amount of prize money in 1961? Use x = 50, as it is 50 years after 1911. Round to the nearest whole number and add three zeros to the end. For instance, if you get an answer like 142.8, this rounds to 143, and add 3 zeros to get $143,000.

OpenStudy (anonymous):

y = 10(1.0605)^x, where x = 50 y = 10(1.0605)^50 = 10(18.86) = 188.6 A. $189,000

OpenStudy (anonymous):

lolol i googled it i love google lol